# From Instruction to Insight: Exploring the Functional and Semantic Roles of Text in Interactive Dashboards

### v1

@@ Introduction
Acknowledging the growing relevance of text in data visualization and the current landscape of generative AI for text, our work aims to understand and formalize the use of text in dashboards, and to provide guidance for dashboard authors and dashboard tool designers alike. Our approach was informed by (a) a survey and qualitative analysis of 190 publicly available dashboards to catalog features of text and characterize their impact on dashboard design and communication, and (b) design review sessions with 13 dashboard creators to reflect and learn from on their personal practices for text use (§2.2). Our contributions are threefold. First, we (i) outline current practices for text in dashboards based on dashboard analysis and expert feedback (§3), which helped inform a typology of text elements in dashboards and a compilation of text use patterns. Second, we discuss (ii) recommended practices for text use in dashboards, based on an adaptation on Setlur et al.’s dashboard heuristics [48] and validated by experts (§4). Third, we outline (iii) future opportunities for the usage of text in dashboards, reflecting on emerging challenges for text use and potential solutions based on the state of the art in data visualization research (§5). Our work sheds light onto the functional, structural, and semantic aspects of text and reflects on its vital role in dashboards for analytical and navigation support. Our investigation shows that text is much more than a mere accompaniment to visuals but rather a fundamental enabler for narrative, explanation, and insight discovery.
@@ 3.1 Text Components
     Interaction Guidance  (58%). Text that describes possible interactions with the dashboard, often written in imperative language, e.g., “click here to filter" (Fig. 1(J)).
@@ 3.2 Semantic Levels
  • Level 3 (LV3): Perceptual and cognitive phenomena (27%). Refers to more complex data statements, such as higher-level trends and pattern synthesis, while using more “natural-sounding" language than LV2 statements. Most often found in  content blocks , and occasionally  chart annotations  and  subheadings .
@@ 3.3.1 Dashboard Text Practices are Based on Experience
When asked about their text use more generally, most participants (8/13) stated that they are not explicitly formalized ($P1$, $P4$, $P6$, $P8$) nor taught ($P6$), but rather based on experience, exposure, and honed senses, e.g., learning from others ($P6$, $P7$, $P13$), learning on the job ($P5$, $P12$), and developing a feel for “what looks good” to them ($P3$, $P7$), which feeds into an implicit set of shared practices: ‘even if it hasn’t been formalized, there is kind of this collective understanding (..) because you’ve seen it repeated in other forms of conception” ($P4$). The more explicit guidance emerging from expert sessions was on structure and formatting guidelines borrowed from other established fields, such as publishing & typography (e.g., magazines) ($P1$, $P4$, $P7$), and graphic & web design ($P5$, $P7$, $P13$).
@@ 3.3.4 Text Formatting for Analytical Support and Emphasis
While seemingly obvious, these text practices require deliberate action from dashboard designers. There is also general consensus that text formatting affordances in dashboard tools are minimal, and scriptable tasks like aligning text color to match visual marks require much manual intervention. Unsurprisingly, several experts reported the use of style guides in their corporate practice ($P4$, $P7$, $P11$), to help inform and align formatting practices across larger design teams.
@@ 3.3.5 Managing Content Overload with Details on Demand
Regarding overall dashboard content, many experts expressed a desire to minimize “big walls of text" ($P1$, $P11$, $P12$), following a general understanding that users “tend to glance over” text content ($P1$, $P4$, $P8$, $P13$): “my approach is to shorten everything that you can shorten” ($P11$). These considerations are particularly relevant in the context of explanatory dashboards, where experts report being very deliberate in choosing what content to include and trying to “keep it as relevant as possible” ($P7$) by focusing on the key takeaways ($P4$, $P6$, $P7$): “I’m constantly questioning myself. Why am I not including this? what is the ulterior motive of why you’re doing this?” ($P4$).
@@ 3.3.7 Dynamic Text for Analytical Support and Breadcrumbs
Less common but interesting uses of dynamic text are on dynamic takeaways, i.e., pre-computed text findings that change depending on the filter selections (e.g., (#53) (#149) (#190)); and a circular “news ticker" strips with data highlights (e.g., (#22) (#189)). On the latter, some experts commented on the value of animated features for user engagement ($P3$, $P7$, $P9$, $P10$): “Maybe you hover on top of a tiny text and it’s just goes bigger, little things that just keep you engaged” ($P7$); and some also mentioned practical uses of dynamic formatting and alerts in their corporate practice ($P4$, $P6$): “if it’s an overall data issue, it’s at the very top and it gets color-coded by brand” ($P4$). But despite the excitement and the meaningful use cases for dynamic text and dynamic formatting, adding such dynamic behaviors with mainstream dashboard tools is, again, reportedly challenging ($P2$, $P5$, $P6$, $P13$).
@@ 4.1 Heuristics for Use of Text in Dashboards
Our construction process started by assessing each of the 39 heuristics in Setlur et al. [48] (H1-H39) and transforming relevant ones to be centered around text use. Some heuristics made explicit mentions of the appropriate use of text (e.g., H6 “the text in the dashboard is legible, easy to read, and useful", here mapping to HP1) and are used almost (if not) verbatim. Others covered communication aspects where text plays an implied or supporting role, which we then redrafted to explicitly mention text as an actor, e.g., H15Charts within the dashboard are at the appropriate level of detail to convey the intended message.”, here mapping to HS2, “Text content within the dashboard are at the appropriate level of detail to convey the intended message”. Heuristics with similar meaning after these redraftings were merged (e.g., H9 thru H12, now HP3). While we iteratively fine-tuned descriptions and rubrics, the essence and scope of the heuristics remained.
@@ 4.3 Expert Feedback on Heuristics
Expert feedback on heuristics was largely positive. The empirical nature of text practices is again acknowledged ($P4$, $P6$, $P12$): e.g., “I teach data visualization frequently (…) [and] we don’t really talk about text that much, or how to use text appropriately” ($P6$); and that having these practices formalized as heuristics is useful (10/13): “For those who are not familiar with the norms or may come from an experience where the norms are different, [this formalization] is a unifier ($P4$). Cited benefits and uses entailed: fostering intentional and “thoughtful usage of text" ($P5$, $P6$, $P13$); guidance for novice designers ($P6$, $P8$) and sanity checks for experienced designers ($P5$, $P9$); and providing a base for more bespoke guidance in applied scenarios ($P7$). Participants also stated that our set of heuristics felt complete (8/13) and representative of their practices (5/13), and they also appreciated the structure with four guidance categories ($P3$, $P4$, $P7$).
While the heuristics were found to provide “a good template” ($P7$) for recommended text practices, many participants argued more specificity is needed to support practical hands-on support ($P3$, $P4$, $P8$, $P9$), e.g., “If it’s just guidance to point people in the right direction, great. But if you want people to have a standard and follow, I need a little more functional direction and just something to grasp and hold on to” ($P4$). A participant also remarked that the heuristics on their own will not ensure dashboard quality, e.g., “you tick the boxes altogether and then maybe you might get an ideal dashboard, but sometimes you might not” ($P12$). Topic suggestions for more specific guidance entailed best practices in text formatting ($P4$, $P5$, $P7$), guidance for annotations ($P3$), and an example catalog to showcase both good and bad applications of individual heuristics ($P3$, $P5$, $P11$, $P13$). More generally, some also underscored the importance of “context and perspective of the user.” ($P10$, $P12$), and better communicate the extent to which each heuristic matters on different use cases ($P10$).
@@ 7 Conclusion
Our work highlights the fundamental role of text in the context of dashboard design and communication, wherein textual elements do much more than accompany visualizations; they enrich and clarify, providing essential context and direction. We hope this paper provides useful insights and future directions for not only dashboard authoring tools to better leverage the utility of text with visualizations, but also to guide effective analytical conversation and text generation in dashboard-adjacent text-data spaces, such as data stories [50, 49], data videos [70], computational notebooks [32], visualization thumbnails [21], and chart captioning [33, 54]. This synergy between text and visuals reaffirms the adage, “A picture is worth a thousand words, but a word can paint a thousand pictures,” underscoring the dual power of text to distill complex data into information understanding.

### v2

@@ Introduction
Acknowledging the growing relevance of text in data visualization and the current landscape of generative AI for text, our work aims to understand and formalize the use of text in dashboards, and to provide guidance for dashboard authors and dashboard tool designers alike. Our approach was informed by (a) a survey and qualitative analysis of 190 publicly available dashboards to catalog features of text and characterize their impact on dashboard design and communication, and (b) design review sessions with 13 dashboard creators to reflect and learn from on their personal practices for text use (§2.2). Our contributions are threefold. First, we (i) outline current practices for text in dashboards based on dashboard analysis and expert feedback (§3), which helped inform a typology of text elements in dashboards and a compilation of text use patterns. Second, we discuss (ii) recommended practices for text use in dashboards, based on an adaptation of Setlur et al.’s dashboard heuristics [48] and validated by experts (§4). Third, we outline (iii) future opportunities for the usage of text in dashboards, reflecting on emerging challenges for text use and potential solutions based on the state of the art in data visualization research (§5). Our work sheds light onto the functional, structural, and semantic aspects of text and reflects on its vital role in dashboards for analytical and navigation support. Our investigation shows that text is much more than a mere accompaniment to visuals but rather a fundamental enabler for narrative, explanation, and insight discovery.
@@ 2.3 Expert Interviews
Callout to text snippets in a dashboard featuring four levels of semantic detail. Level 1:
This is a Gender & Population graph. It shows the percentage of genders in New York City, and changes respectively to the chosen city. By hovering over the gender you will be able to see the total gender count. Level 2: #1 Population In America with 17M Population. Level 3: With 17M total population, 1 in every 38 people in the United States call New York City home. Level 4: For students that wants to live in New York City affordably while exploring the city, it is difficult to find a place to stay.
@@ 3.1 Text Components
     Interaction Guidance  (58%). Text that describes possible interactions with the dashboard, often written in imperative language, e.g., “click here to filter” (Fig. 1(J)).
@@ 3.2 Semantic Levels
  • Level 3 (LV3): Perceptual and cognitive phenomena (27%). Refers to more complex data statements, such as higher-level trends and pattern synthesis, while using more “natural-sounding” language than LV2 statements. Most often found in  content blocks , and occasionally  chart annotations  and  subheadings .
@@ 3.3.1 Dashboard Text Practices are Based on Experience
When asked about their text use more generally, most participants (8/13) stated that they are not explicitly formalized ($P1$, $P4$, $P6$, $P8$) nor taught ($P6$), but rather based on experience, exposure, and honed senses, e.g., learning from others ($P6$, $P7$, $P13$), learning on the job ($P5$, $P12$), and developing a feel for “what looks good” to them ($P3$, $P7$), which feeds into an implicit set of shared practices: “even if it hasn’t been formalized, there is kind of this collective understanding (..) because you’ve seen it repeated in other forms of conception” ($P4$). The more explicit guidance emerging from expert sessions was on structure and formatting guidelines borrowed from other established fields, such as publishing & typography (e.g., magazines) ($P1$, $P4$, $P7$), and graphic & web design ($P5$, $P7$, $P13$).
@@ 3.3.2 Text Components and Semantic Levels
Level 2 content in a dashboard tooltip: "Perception (2) (Harvey Walters) is found in 52% of Amanda Sharpe decks (51 out of 97)". Level 3 content in the same tooltip: "That is 0.04 decks per day, or 1 deck every 25.6 days, based on the most recent of either the card release date (Sep ’20) or investigator release date (Sep ’20)".
@@ 3.3.4 Text Formatting for Analytical Support and Emphasis
While seemingly obvious, these text practices require deliberate action from dashboard designers. There is also a general consensus that text formatting affordances in dashboard tools are minimal, and scriptable tasks like aligning text color to match visual marks require much manual intervention. Unsurprisingly, several experts reported the use of style guides in their corporate practice ($P4$, $P7$, $P11$) to help inform and align formatting practices across larger design teams.
@@ 3.3.5 Managing Content Overload with Details on Demand
Regarding overall dashboard content, many experts expressed a desire to minimize “big walls of text” ($P1$, $P11$, $P12$), following a general understanding that users “tend to glance over” text content ($P1$, $P4$, $P8$, $P13$): “my approach is to shorten everything that you can shorten” ($P11$). These considerations are particularly relevant in the context of explanatory dashboards, where experts report being very deliberate in choosing what content to include and trying to “keep it as relevant as possible” ($P7$) by focusing on the key takeaways ($P4$, $P6$, $P7$): “I’m constantly questioning myself. Why am I not including this? what is the ulterior motive of why you’re doing this?” ($P4$).
@@
Dashboard overlay with instruction callouts over dashboard elements. Examples: (a) "Overall performance: click to switch monitoring metrics". (b) "Update field refreshed daily at 12PM CT".
@@ 3.3.7 Dynamic Text for Analytical Support and Breadcrumbs
Less common but interesting uses of dynamic text are on dynamic takeaways, i.e., pre-computed text findings that change depending on the filter selections (e.g., (#53) (#149) (#190)); and a circular “news ticker” strips with data highlights (e.g., (#22) (#189)). On the latter, some experts commented on the value of animated features for user engagement ($P3$, $P7$, $P9$, $P10$): “Maybe you hover on top of a tiny text and it’s just goes bigger, little things that just keep you engaged” ($P7$); and some also mentioned practical uses of dynamic formatting and alerts in their corporate practice ($P4$, $P6$): “if it’s an overall data issue, it’s at the very top and it gets color-coded by brand” ($P4$). But despite the excitement and the meaningful use cases for dynamic text and dynamic formatting, adding such dynamic behaviors with mainstream dashboard tools is, again, reportedly challenging ($P2$, $P5$, $P6$, $P13$).
@@ 4.1 Heuristics for Use of Text in Dashboards
Our construction process started by assessing each of the 39 heuristics in Setlur et al. [48] (H1-H39) and transforming relevant ones to be centered around text use. Some heuristics made explicit mentions of the appropriate use of text (e.g., H6the text in the dashboard is legible, easy to read, and useful”, here mapping to HP1) and are used almost (if not) verbatim. Others covered communication aspects where text plays an implied or supporting role, which we then redrafted to explicitly mention text as an actor, e.g., H15Charts within the dashboard are at the appropriate level of detail to convey the intended message”, here mapping to HS2, “Text content within the dashboard are at the appropriate level of detail to convey the intended message”. Heuristics with similar meaning after these redraftings were merged (e.g., H9 thru H12, now HP3). While we iteratively fine-tuned descriptions and rubrics, the essence and scope of the heuristics remained.
@@ 4.3 Expert Feedback on Heuristics
Expert feedback on heuristics was largely positive. The empirical nature of text practices is again acknowledged ($P4$, $P6$, $P12$): e.g., “I teach data visualization frequently (…) [and] we don’t really talk about text that much, or how to use text appropriately” ($P6$); and that having these practices formalized as heuristics is useful (10/13): “For those who are not familiar with the norms or may come from an experience where the norms are different, [this formalization] is a unifier” ($P4$). Cited benefits and uses entailed: fostering intentional and “thoughtful usage of text” ($P5$, $P6$, $P13$); guidance for novice designers ($P6$, $P8$) and sanity checks for experienced designers ($P5$, $P9$); and providing a base for more bespoke guidance in applied scenarios ($P7$). Participants also stated that our set of heuristics felt complete (8/13) and representative of their practices (5/13), and they also appreciated the structure with four guidance categories ($P3$, $P4$, $P7$).
While the heuristics were found to provide “a good template” ($P7$) for recommended text practices, many participants argued more specificity is needed to support practical hands-on support ($P3$, $P4$, $P8$, $P9$), e.g., “If it’s just guidance to point people in the right direction, great. But if you want people to have a standard and follow, I need a little more functional direction and just something to grasp and hold on to” ($P4$). A participant also remarked that the heuristics on their own will not ensure dashboard quality, e.g., “you tick the boxes altogether and then maybe you might get an ideal dashboard, but sometimes you might not” ($P12$). Topic suggestions for more specific guidance entailed best practices in text formatting ($P4$, $P5$, $P7$), guidance for annotations ($P3$), and an example catalog to showcase both good and bad applications of individual heuristics ($P3$, $P5$, $P11$, $P13$). More generally, some also underscored the importance of “context and perspective of the user” ($P10$, $P12$), and better communicate the extent to which each heuristic matters on different use cases ($P10$).
@@ 7 Conclusion
Our work highlights the fundamental role of text in the context of dashboard design and communication, wherein textual elements do much more than accompany visualizations; they enrich and clarify, providing essential context and direction. We hope this paper provides useful insights and future directions for not only dashboard authoring tools to better leverage the utility of text with visualizations, but also to guide effective analytical conversation and text generation in dashboard-adjacent text-data spaces, such as data stories [50, 49], data videos [70], computational notebooks [32], visualization thumbnails [21], and chart captioning [33, 54]. This synergy between text and visuals reaffirms the adage, “A picture is worth a thousand words, but a word can paint a thousand pictures”, underscoring the dual power of text to distill complex data into information understanding.